A(n) 90-g ice cube at 0°C is placed in 680 g of water at 26°C. What is the final temperature of the mixture?
Answer #1

Using

Heat lost = Heat gained

Let the final temperature be T

For converting 90g ice to water at 0 deg

H = mCf = 90*333 = 29970 J

Now

29970 + mCw(dT) = MC(dT)

29970 + 90*4.186(T-0) = 680*4.186(26-T)

29970 + 376.74T = 74008.48-2846.48T

3223.22T = 44038.48

T = 13.66o C